Circular No. 5997

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

OPTICAL GRAVTATIONAL LENSING EXPERIMENT (OGLE) B. Paczynski, Princeton University Observatory, writes: “During June

7–Sept. 15, ∼ 2 million stars in the Galactic Bulge will be monitored by the OGLE, and we expect that 2–4 microlensing events, each lasting a few weeks, will be discovered. The new computer system automatically signals the events while they are on the rise, making it possible to carry out photometric and/or spectroscopic follow-up observations. If you would like to be notified about on-going events, please send your request to A. Udalski (udalski@sirius.astrouw.edu.pl). More information can be obtained over Internet via ‘sirius.astrouw.edu.pl’ by anonymous ftp (directory ‘/ogle’, file ‘early.warning’).”

1993 FW J. Davies, Joint Astronomy Centre, Hilo, reports that a search for 1993

FW on Apr. 27.35 UT with the new infrared camera IRCAM3 at the U.K. Infrared Telescope did not detect the object (ephemeris on IAUC 5939): “In a 2′′ software aperture, the deduced 5σ upper limit at J is 22.0. Combining this with an estimated V magnitude of 23.3 implies a V −J color of < +1.5, less extreme than both (5145) Pholus (V − J = +2.46; Davies et al. 1993, Icarus 102, 166) and 1993 HA_{2} (V − J = +2.0 ± 0.2; unpublished). Since this implies that some Kuiper-belt objects may be less red than Pholus and 1993 HA_{2}, further observations are urgently required.”

COMET TAKAMIZAWA-LEVY (1994f)

Extension to ephemeris on IAUC 5985:

1994TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
June 1 17^{h}06.^{m}40 +69^{o}27.^{′}6 0.945 1.366 88.^{o}4 47.^{o}9 8.2
3 16 18.1 +70 13.0 0.968 1.370 87.4 47.7 8.3
5 15 30.8 +70 06.5 0.994 1.374 86.3 47.5 8.4
7 14 48.29 +69 16.6 1.025 1.378 85.0 47.2 8.4
9 14 12.43 +67 55.4 1.059 1.383 83.6 46.8 8.5
11 13 43.27 +66 14.7 1.097 1.389 82.1 46.4 8.6
13 13 19.91 +64 23.5 1.137 1.396 80.6 45.9 8.7
15 13 01.25 +62 28.2 1.180 1.402 79.0 45.3 8.8
17 12 46.26 +60 32.9 1.225 1.410 77.4 44.7 8.9

1994 May 25 Daniel W. E. Green